Manish travels a certain distance by his car at the rate of 12km/hr and walks back at the rate of 3km/hr . the whole journey took 5 hrs . what is the distance he covered on car

Answer by stanbon(75887)

Manish travels a certain distance by his car at the rate of 12km/hr and walks back at the rate of 3km/hr . the whole journey took 5 hrs . what is the distance he covered on car
------
Car DATA:
rate = 12 km/hr ; time = t hrs : distance = 12t km
------
Walking DATA:
rate = 3 km/hr ; time = 5-t hrs ; distance = 3(5-t) km
-----
Equation:
car dist = walk dist
12t = 15-3t
15t = 15
time = 1 hr
------
Distance covered by car: 12t = 12 km
=========================================
Cheers,
Stan H.
===================